OpenStudy (supercait10):

Question 1. What is the solution to 8/9 =-7 ? A. 81 B. 36 C. -36 D. -81 Question 2. Which Expression can be written as 9 . (12+7)? A. 9 + (21+7) B. (9+12) + (9+7) C. 9 . 12+7 D. (9.12) + (9.7) Question 3. Which property is show by 2+5=5+2? A. Associative Property of Addition B. Commutative Property of Addition (picked this one) C. Distributive Property D. Identity Property of Addition Question 4. How is the sentence "Six more than x is nine) written as an equation? A. 6=x+9 B. 6+9=x C 6 + x= 9 D. 6+9+x=0 Question 5.Which Property is shown by 4 +(5+6)=(4+5) +6 A. Commutative Property of Addition B. Distributive Property C. Additive Identify (picked this one) D. Associative Property of Addition Question 6. Which symbol will make |-8| ? |-10| A. > (picked this one) B. < C. = Question 6. What is the solution to Q/-7=6 A. Q=-42 B. Q=-13 C.Q=13 Q=42 Just need help like explaining not given the answers

OpenStudy (supercait10):

\[\frac{ B }{ 9 } = -7\]

OpenStudy (kittiwitti1):

Problem 1.\[\frac{B}{9}=-7\rightarrow B=-7\times9=?\] Problem 2. I'm not sure what you mean by the dot in 9.(12+7), but if it means a dot like •, that means multiply. Then you would get \[9\times(12+7)=9(12+7)=\text{ by distribution }(9\times12)+(9\times7)\] Problem 3. You are right because the Commutative Property is the one that moves numbers around with the same numerical result:\[\text{An example problem: }a+b=b+a\] Problem 4. "Six more than x is nine"\[\text{"Six more than x" means add 6 to x }x+6\]\[\text{Since we know that equals nine you put = 9 at the end }\]\[x+6=9\] Problem 5. Wrong, the Additive Identity is "add 0 to a number and you will get the same exact number". The answer is Associative Property of Addition, because you can regroup the numbers and still get the same result.\[\text{Example problem: }(a+b)+c=a+(b+c)\] Problem 6. An absolute value is the distance the integer is from 0 on a number line. (If you still don't get it just ask) Problem "6" (7). Isolate the Q and get the answer:\[\frac{Q}{-7}=6\rightarrow Q=-7\times6=?\]

OpenStudy (kittiwitti1):

Problem 4 typo:\[\text{"Six more than x" means add 6 to x }\rightarrow x+6\] If you got confused on Problem 6:\[\text{example: }\left|-a\right|=a\]

OpenStudy (kittiwitti1):

http://prntscr.com/8g573w If you can't see the stuff I typed, click this link.

OpenStudy (kittiwitti1):

ADDITIONAL EXPLANATIONS: Problems 1 and 7: When faced with a single variable in an equation, and asked to simplify/solve/evaluate, isolate the variable and simplify the other side of the equation. Problem 2: Distributive Property is a(b + c)= ab + ac and so on.

OpenStudy (zzr0ck3r):

where are you all seeing a B in problem 1?

OpenStudy (kittiwitti1):

@zzr0ck3r She corrected the 8 to a B in her first comment.

OpenStudy (zzr0ck3r):

weird because -7*9=-63 is not an option

OpenStudy (supercait10):

that is what the equation says I'm going off by what it says just don't give me the answers just explain it

OpenStudy (kittiwitti1):

Well, technically... by explaining, we're also giving you the answers, but allowing you to understand and learn at the same time. :] Because giving direct answers is against the OpenStudy Code of Conduct. ^^